OpenStudy (anonymous):

Use substitution to solve. 5x-y=10 7x-2y=11

OpenStudy (jim766):

to solve a system using substitution, you need to have at least one of the variables alone 5x - y = 10 + y + y 5x = 10 + y -10 -10 5x-10 = y do you see how I did that? 7x -2y = 11